OpenStudy (anonymous):

What is the solution to the equation 9^3x ≈ 7 ? x = 0.376 x = 0.295 x = –0.295 x = –0.376 What is the solution to the equation 9^3x ≈ 7 ? x = 0.376 x = 0.295 x = –0.295 x = –0.376 @Mathematics

OpenStudy (anonymous):

3x ln(9)= ln(7) 3x = ln(7)/ln(9) x= 1/3 (ln(7)/ln(9))

OpenStudy (anonymous):

x = 0.295